OpenStudy (anonymous):

16 3/4 % write in decimals....

OpenStudy (anonymous):

is it like \[16\frac{ 3 }{ 4 }%\]%

OpenStudy (insa):

16.75 %

OpenStudy (anonymous):

Mutiply 16.75 by four and put it over 400 and simplify

OpenStudy (anonymous):

yes it \[16\frac{ 3 }{ 4}\]

OpenStudy (anonymous):

so 16.75% and in decimal 16.75/100 =.1675